please help....
2.2x-5=2(1.4x+3)

Answer by RAY100(1637)

2.2x -5 = 2(1.4x +3)
.
2.2x -5 = 2.8x +6
.
-11 = .6x
.
x = - 18 1/3
.
check
.
2.2(-18.33) -5 = 2( 1.4 (-18.33) +3)
.
-45.333 = -45 .333,,,,ok